OpenStudy (anonymous):

ten times the square of a nonzero number is equal to eitghty times the number. what is the number?

OpenStudy (justin_lee):

Can you write an equation out of this? It will help you.

OpenStudy (justin_lee):

^ Yup.

OpenStudy (abhisar):

Hey there! Let the non-zero number be \(\sf x\), then according to the question, \(\sf 10 \times x^2 = 80 \times x\) Now solve the equation to find the value of \(\sf x\)

OpenStudy (solomonzelman):

good set up, cookielate. BUT by dividing by x, you lose solutions

OpenStudy (solomonzelman):

well, if your condition is that \(\large\color{black}{ \displaystyle x>\bf 0 }\), but I didn't see that.

OpenStudy (solomonzelman):

my replies got mixed up-:(

OpenStudy (anonymous):

8 was correct. you guys have no idea ho much yall helped thank you so much

OpenStudy (solomonzelman):

anyway, \(\large\color{black}{ \displaystyle10x^2 = 80x }\) \(\large\color{black}{ \displaystyle10x^2 - 80x=0 }\) \(\large\color{black}{ \displaystyle x^2 - 8x=0 }\) \(\large\color{black}{ \displaystyle x(x-8)=0 }\) yes, ultimately you come to one answer only though, as it says that x is a none zero number. and the constant "c" is not there, so you took a "shortcut" cutting it out, knowing that 0 will be a solution of a quadratic.

OpenStudy (solomonzelman):

I was going to medal that reply, because it is eventually correct... and youerased it... dky

OpenStudy (abhisar):

How about this? \(\huge \tt 10x^2 =80x \\ \huge \frac{10x^2}{x}=80 \\ \huge 10x=80 \\ \huge \Rightarrow x=8\)

OpenStudy (solomonzelman):

I am perfectly fine with that! you can divide by x, yes, because the solution you are losing is 0, but the problem already told us that \(\large\color{black}{ \displaystyle x \ne 0}\)

OpenStudy (solomonzelman):

(but, in other cases don't cancel out the x like this)